# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 1 Introduction
One relatively simple way to evaluate bootstrap current in stellarators is to use the long mean free path asymptotic formula of Shaing and Callen (1983) which contains all the information about device geometry in a geometrical factor independent of plasma parameters. This way is especially suited for stellarator optimization (Beidler et al., 1990; Nakajima et al., 1989) where multiple fast estimates of bootstrap current are required. Despite that it has been derived more than 40 years ago, the validity range of this formula still remains unclear. Although the same result is exactly (Helander et al., 2011) or approximately (Boozer and Gardner, 1990) reproduced by different derivations, however, it is not reproduced in the $1/\nu$ regime by numerical modelling (Beidler et al., 2011). Namely, with decreasing normalized collisionality $\nu_{\ast}$, the bootstrap coefficient does not come to a saturation at the asymptotic limit but rather shows a complicated behavior which is quite different at magnetic surfaces with different radii (Kernbichler et al., 2016). In turn, in the presence of radial electric field, numerical modelling shows a saturation to the Shaing-Callen limit with decreasing $\nu_{\ast}$. This feature of the bootstrap coefficient regularly observed in DKES (Hirshman et al., 1986; van Rij and Hirshman, 1989) modelling, and later confirmed by various codes of different types (Beidler et al., 2011) has been pointed out to the authors of this paper by Henning Maassberg (2004) more than 20 years ago, and, together with the problem of bootstrap resonances (Boozer and Gardner, 1990), it was the main reason to start the development of the drift kinetic equation solver NEO-2 (Kernbichler et al., 2008; Kasilov et al., 2014; Martitsch et al., 2014; Kernbichler et al., 2016; Kapper et al., 2016, 2018). Finally, we can present here the results of this long going effort.
@@
We restrict our analysis here to the mono-energetic approach (Beidler et al., 2011) employing the Lorentz collision model, since this approach is sufficient for the account of main effects related to the magnetic field geometry. Effects of energy and momentum conservation pertinent to the full linearized collision model can be recovered then with good accuracy from the mono-energetic solutions of the kinetic equation using various momentum correction techniques (Taguchi, 1992; Sugama and Nishimura, 2002, 2008; Maassberg et al., 2009). This mono-energetic approach is briefly outlined in Section 2.1 where also the basic notation is introduced. In Section 2.2 we re-derive the Shaing-Callen formula (Shaing and Callen, 1983) within this approach in the $1/\nu$ transport regime in order to obtain an explicit solution for trapped particle distribution accounting for various trapped particle classes. The structure of this solution is quite demonstrative for the reasons behind the “anomalous” behavior of bootstrap current mentioned above. In that section, we also extend the alternative derivation of Helander et al. (2011) within the adjoint approach to the next order in collisionality. Although the obtained correction has no effect on the resulting Ware pinch (bootstrap) coefficient, it is useful for understanding a certain paradox contained in this solution. Collisionless asymptotic solutions are compared to numerical solutions for finite plasma collisionality in Section 2.3 where the cases with convergence of these solutions to the Shaing-Callen limit at low collisionality are demonstrated. In Section 3 we examine the effect of collisional boundary layers on the distribution function in the adjoint (Ware pinch) problem where they lead to the off-set of this function from the value of collisionless asymptotic of Helander et al. (2011) and respective off-set of the Ware-pinch coefficient. In particular, in Section 3.2 we introduce a simplified approach (propagator method) to describe this off-set in the leading order over collisionality with help of a set of Wiener-Hopf type integral equations. In Section 3.3, with the help of this set, we derive the conditions on the equilibrium magnetic field required to avoid the leading order off-set. We obtain a simple expression for the distribution function off-set in case these conditions are weakly violated in Section 3.5. In Section 4 we examine the off-set of the distribution function and Ware pinch coefficient at irrational flux surfaces both, numerically, with help of NEO-2 code solutions at high order rational magnetic field lines approximating the irrational surface (Section 4.1), and using the analytical estimates of the asymptotic behavior of the off-set with reducing plasma collisionality (Section 4.2). A related issue of bootstrap resonances is briefly discussed in Section 4.4. In Section 5 we study the effect of banana orbit precession (in particular, due to a finite radial electric field) on the off-set of Ware pinch coefficient and formulate there a simple bounce averaged approach for the account of this effect in computations of Ware pinch coefficient using the numerical solutions for the distribution function in the $1/\nu$ regime. A qualitative discussion of the off-set in the direct problem describing bootstrap coefficient is presented in Section 6 where we also examine the possibility of bootstrap effect at the magnetic axis. Finally, the results are summarized in Section 7 where some implications for reactor optimization are also discussed.
@@ 2.1 Adjoint mono-energetic problems on a closed field line
For the present analysis of bootstrap current convergence with plasma collisionality, a mono-energetic approach is sufficient where a Lorentz collision model and constant electrostatic potential within flux surfaces are assumed. The linear deviation of the distribution $f$ function from the local Maxwellian $f_{M}$ is presented as a superposition of thermodynamic forces $A_{k}$ as detailed in (Kernbichler et al., 2016),
@@
$H_{\varphi}^{\prime}=\frac{|\lambda|}{3B^{\varphi}}\left(3+\lambda^{2}\right)|\nabla r|k_{G}\rho_{L}.$ (5)
Here, we use a field aligned coordinate system $(r,\vartheta_{0},\varphi)$ where $r$ is a flux surface label (effective radius) fixed by the condition $\langle|\nabla r|\rangle=1$, $\vartheta_{0}=\vartheta-\iota\varphi$ is a field line label, $\iota$ is the rotational transform and $\vartheta$ and $\varphi$ are poloidal and toroidal angle of periodic Boozer coordinates (Boozer, 1981; D’haeseleer et al., 1991), respectively. Variables in the velocity space are parallel velocity sign $\sigma=v_{\parallel}/|v_{\parallel}|$ and two invariants of motion, $z=m_{\alpha}v^{2}/(2T_{\alpha})$ and $\eta=v_{\perp}^{2}/(v^{2}B)$ respectively being the normalized kinetic energy (playing a role of parameter) and perpendicular adiabatic invariant (magnetic moment). The other notation is pitch parameter $\lambda=v_{\parallel}/v=\sigma\sqrt{1-\eta B}$, the mean free path $l_{c}=v/(2\nu_{\perp})$ defined via deflection frequency $\nu_{\perp}$ (the same as $\nu$ in Ref. (Beidler et al., 2011)), radial guiding center velocity $v_{g}^{r}={\bf v}_{g}\cdot\nabla r$, Larmor radius $\rho_{L}=cm_{\alpha}v/(e_{\alpha}B)$ and the geodesic curvature given by
@@
with ${\bf h}={\bf B}/B$ and $\psi$ being the normalized toroidal flux, and $B^{\varphi}$, $B_{\varphi}$ and $B_{\vartheta}$ being contra- and covariant components of the magnetic field in Boozer coordinates.
@@
For the present analysis, we solve the kinetic equation on the rational surface using a “representative” field line where necessary conditions valid for the irrational flux surface are fulfilled. Flux surface average (10) at the irrational surface corresponds then to the limit of the series of representative field lines closed after $N$ turns at respective rational surfaces, $r=r_{N}=M/N$, which converge to the irrational surface, $\lim\limits_{N\rightarrow\infty}\iota(r_{N})=\iota(r)$. The requested condition is Liouville’s theorem, which states that the integral over any closed surface of the normal component of the guiding center velocity multiplied by the phase space Jacobian is zero for fixed total energy and perpendicular adiabatic invariant used as phase space variables. For the magnetic surface with constant electrostatic potential, this means $\left\langle Bv_{g}^{r}/v_{\parallel}\right\rangle=0$, where the surface integration is performed over regions where $v_{\parallel}^{2}=v^{2}(1-\eta B)\geq 0$ keeping invariants $z$ and $\eta$ constant. The field line average form (10) of Liouville’s theorem results in
@@
for passing and trapped particles, respectively (see (4)). We will call these conditions “quasi-Liouville’s theorem”. In the devices with stellarator symmetry conditions (15) are satisfied for closed field lines passing through the magnetic field symmetry point what is obvious due to anti-symmetry of the geodesic curvature (and, respectively, of $v_{g}^{r}$ and $s_{(1)}$) with respect to this point, $s_{(1)}(2\varphi_{0}-\varphi)=-s_{(1)}(\varphi)$. Restricting our analysis to such devices with a single global maximum per field period, the reference field line is then the one starting from the global maximum, which must be located at the symmetry point.
@@ 2.2 Collisionless asymptotic solutions
where $\varphi_{\text{beg}}=\varphi_{j}^{-}(\eta)$ for trapped particles and $\varphi_{\text{beg}}=\varphi_{0}$ for passing, and we require that each of $g_{n}$ is continuous at the periodic boundary and at the turning points. Continuity at $\varphi=\varphi_{j}^{-}$ means that $\bar{g}_{n}$ is an even function of $\sigma$ in the trapped region, $\bar{g}_{n}=\bar{g}_{n}(\eta)$, while continuity at the periodic boundary in the passing region and at $\varphi=\varphi_{j}^{+}$ in the trapped region results in equation for the integration constant $\bar{g}_{n-1}$ (solubility constraint for $g_{n}$). The leading order constraint for $g_{0}$ gives a bounce-averaged equation for $g_{-1}$
@@
where $\varphi_{\text{end}}=\varphi_{j}^{+}(\eta)$ for trapped particles and $\varphi_{\text{beg}}=\varphi_{N}$ for passing, and the sum $\sum\limits_{\sigma=\pm 1}^{\text{trapped}}$ is taken for trapped particles only. Higher order constraints result in equations for integration constants $\bar{g}_{n}$,
@@
For the next order correction $g_{0}$, we notice that first two terms in the r.h.s. of Eq. (19) are odd in $\sigma$. Therefore, equation (23) for the integration constant (which can only be even) is homogeneous resulting in $\bar{g}_{0}=\text{const}$. Due to the continuity of the distribution function this is the same constant as $g_{-1}$ in the passing region where we have already set it to zero. Thus, substituting in (19) the derivative $\partial g_{-1}/\partial\eta$ via (27) and the source $s=s_{(1)}$ via (4) and (5) we get explicitly
@@
where we have exchanged the derivative over $\eta$ with integration in the first term using $H_{\varphi}^{\prime}(\varphi_{j}^{-}(\eta),\eta)=0$. It should be noted now that the integral over $\varphi^{\prime}$ is a discontinuous function of $\eta$ at the boundaries between classes $\eta=\eta_{c}$ where either $\varphi_{j}^{-}(\eta)$ or $\varphi_{j}^{+}(\eta)$ jumps. Respectively, function $g_{0}$ contains a $\delta$-function $\delta(\eta-\eta_{c})$, which is required by particle conservation in the boundary layer. Namely, integrating Eq. (3) over $\eta$ across the boundary we get
@@
$C_{\parallel}=C_{\parallel}(r)=\frac{1}{\rho_{L}}\sum_{\alpha}C_{\alpha}\rho_{L}=\frac{3c}{4B\rho_{L}}\frac{\partial p}{\partial r},\qquad p=\sum_{\alpha}p_{\alpha},$ (38)
and we used the condition $\bar{g}_{0}=0$ at the boundary $\eta=1/B$ when integrating by parts in the last expression (37). According to (35), term $g_{0}-\bar{g}_{0}$ is a derivative which contributes only at the lower limit $\eta=0$ where using explicitly (5) one gets parallel equilibrium current density as
@@
which is a “true” rational surface condition (Solov’ev and Shafranov, 1970). Using the explicit form of $H^{\prime}_{\varphi}$ in (34) and taking the flux “surface average” (10) of $j_{\parallel}B$ thus eliminating the Pfirsch-Schlüter current which is driven solely by charge separation potential and, therefore, must satisfy $\langle j_{\text{PS}}B\rangle=0$, we finally obtain bootstrap current density $j_{b}$ which scales with $B$ as
@@
The separate contribution of trapped particle region to $\lambda_{bB}$ is obtained by replacement of lower integration limit in (37) from 0 to $\eta_{b}+o$,
@@
The factor (42) is the same with the result of Shaing and Callen, what is a natural consequence of using the same method. Namely, $\lambda_{bB}=(1-f_{c})\langle\tilde{G}_{b}\rangle/S$ where $S={\rm d}V/{\rm d}r$ is the flux surface area, fraction of circulating particles $f_{c}$ is given by Shaing and Callen (1983), Eq. (56) and geometrical factor $\tilde{G}_{b}$ is given there by Eqs. (75b) and (60c). In case of negligible toroidal equilibrium current, $B_{\vartheta}=0$, the result of Boozer and Gardner (1990) is also approximately recovered, $\lambda_{bB}=\Delta_{0}B_{\varphi}{\rm d}r/{\rm d}\psi+O(\varepsilon_{\text{M}})$, where the quantity $\Delta_{0}$ is given by Eqs. (51), (53) and (54) of Boozer and Gardner (1990) ignoring small trapped particle contribution and other corrections linear in $\varepsilon_{M}$ (in case of a circular tokamak, $\lambda_{bB}\approx A\Delta_{0}$ where $A$ is the aspect ratio).
@@
Note that the trapped particle distribution function here is formally the same with the one given in the implicit form by Eq. (54) of Shaing and Callen (1983) omitting there the ripple plateau contribution absent in our case where we ignored banana precession (cross-field drift within flux surfaces) already in the starting equation (3). On the other hand, we obtained also the explicit expression for its odd part (35) which is demonstrative for the distribution of parallel equilibrium current in the velocity space. It can be seen that a significant part of this current (essentially of the Pfirsch-Schlüter current) flows in the boundary layers between trapped particle classes (but not in the main, trapped-passing boundary layer). The origin of these localized currents is the same with the origin of Pfirsch-Schlüter current in all devices, i.e. they remove charge separation introduced by the radial particle drift. However, in contrast to tokamaks (and tokamak-like part of the current in stellarators) where compensation currents are produced in a long mean free path regime within a single turn of particle bounce motion due to the finite radial displacement during this time, localized currents serve to compensate charge separation by finite bounce-averaged drift, which accumulates on much longer collisional detrapping time. Interpreting the source term $s_{(1)}$, Eq. (4), in the regions where it is positive as a source of “particles” and where it is negative as a source of “anti-particles” whose total amount generated at the flux surface is the same with the amount of “particles” being a consequence of quasi-Liouville’s theorem (15), we see that particles (or anti-particles) accumulated in the local ripple wells due to the finite bounce averaged drift, $H_{j}\neq 0$, can only leave the wells due to collisional scattering flux through the class boundaries (see Fig. 1). Since local collisional flux density generally needs not to be continuous at this boundary, $\partial g_{-1}(\eta-o)/\partial\eta\neq\partial g_{-1}(\eta+o)/\partial\eta$, a significant amount of particles (anti-particles) is re-distributed through the boundary layers, what is manifested by the $\delta$-like behavior of $g_{0}$, Eq. (35), being, up to a factor, a parallel flow density in velocity space.
@@
The asymptotic series expansion (17) can be continued to the next order, leading to the correction $g_{1}$. This correction, however, has a non-integrable singularity at class transition boundaries, $g_{1}\propto(\eta-\eta_{c})^{-2}$ (in contrast to integrable singularity of $g_{0}\propto\log|\eta-\eta_{c}|$), which is a consequence of finite geodesic curvature at respective local field maxima. Therefore, contributions of the corrections to the collisional flux density in the matching conditions (31) become compatible at finite collisionality for $|\eta-\eta_{c}|=\delta\eta\propto l_{c}^{-1/2}\propto\nu^{1/2}$ where $\partial g_{0}(\eta)/\partial\eta\sim\delta\eta^{-1}$ and $\partial g_{1}(\eta)/\partial\eta\sim l_{c}^{-1}\delta\eta^{-3}$. In other words, the expansion (17) breaks down at the edge of the boundary layer of width $\delta\eta$. Respectively, matching interval $o$ in Eq. (28), which is an infinitesimal for vanishing collisionality, should satisfy $\delta\eta\ll o\ll\eta$ for finite collisionality, which means that the error of the asymptotic solution scales with $\sqrt{\nu}$ rather than with $\nu$. Since particle – anti-particle re-distribution flux manifested by $\delta$-functions in asymptotic expression (35) for $g_{0}$ is independent of collisionality, we can estimate the odd part of the distribution function which carries this flux in the class-transition boundary layers as $g_{\text{odd}}\propto\delta\eta^{-1}\propto l_{c}^{1/2}\propto\nu^{-1/2}$. As shown below, presence of these strong parallel flows in class transition boundary layers is actually responsible for the “anomalous” behavior of bootstrap current at finite collisionality.

### v3

@@ 1 Introduction
One relatively simple way to evaluate bootstrap current in stellarators is to use the long mean free path asymptotic formula of Shaing and Callen (1983) which contains all the information about device geometry in a geometrical factor independent of plasma parameters. This way is especially suited for stellarator optimization (Beidler et al., 1990; Nakajima et al., 1989) where multiple fast estimates of bootstrap current are required. Despite that it has been derived more than 40 years ago, the validity range of this formula still remains unclear. Although the same result is exactly (Helander et al., 2011) or approximately (Boozer and Gardner, 1990) reproduced by different derivations, however, it is not reproduced in the $1/\nu$ regime by numerical modelling (Beidler et al., 2011). Namely, with decreasing normalized collisionality $\nu_{\ast}$, the bootstrap coefficient does not come to a saturation at the asymptotic limit but rather shows a complicated behavior which is quite different at magnetic surfaces with different radii (Kernbichler et al., 2016). In turn, in the presence of radial electric field, numerical modelling shows a saturation to the Shaing-Callen limit with decreasing $\nu_{\ast}$. This feature of the bootstrap coefficient regularly observed in DKES (Hirshman et al., 1986; van Rij and Hirshman, 1989) modelling, and later confirmed by various codes of different types (Beidler et al., 2011) has been pointed out to the authors of this paper by Henning Maassberg (2004) more than 20 years ago, and, together with the problem of bootstrap resonances (Boozer and Gardner, 1990), it was the main reason to start the development of the drift kinetic equation solver NEO-2 (Kernbichler et al., 2008; Kasilov et al., 2014; Martitsch et al., 2016; Kernbichler et al., 2016; Kapper et al., 2016, 2018). Finally, we can present here the results of this long going effort.
@@
As shown below, for the “anomalous” behaviour of the bootstrap coefficient, the interaction between boundary layers separating co- and counter-passing particles from trapped particles and the layers separating different trapped particle classes from each other plays an important role. In collisionless asymptotic theories, these layers are assumed infinitely thin and non-overlapping. This cannot be fulfilled at irrational flux surfaces, where the number of trapped particle classes is infinite, while the width of the layers is finite at any collisionality. Another reason is the interaction of the trapped-passing boundary layer with itself, which happens in case of anti-sigma configurations where the magnetic field maximum at a given flux surface is reached on a line which splits the field line into non-equivalent segments exchanging transient particles via the boundary layer. In both cases, an important prerequisite for anomaly is a finite bounce-averaged cross-surface drift of trapped particles, which is absent in axisymmetric and nearly absent in sufficiently accurate quasi-symmetric configurations showing no anomalies of the bootstrap current (Landreman et al., 2022).
@@
We restrict our analysis here to the mono-energetic approach (Beidler et al., 2011) employing the Lorentz collision model, since this approach is sufficient for the account of main effects related to the magnetic field geometry. Effects of energy and momentum conservation pertinent to the full linearized collision model can be recovered then with good accuracy from the mono-energetic solutions of the kinetic equation using various momentum correction techniques (Taguchi, 1992; Sugama and Nishimura, 2002, 2008; Maassberg et al., 2009). This mono-energetic approach is briefly outlined in Section 2.1 where also the basic notation is introduced. In Section 2.2 we re-derive the Shaing-Callen formula (Shaing and Callen, 1983) within this approach in the $1/\nu$ transport regime in order to obtain an explicit solution for trapped particle distribution accounting for various trapped particle classes. The structure of this solution is quite demonstrative for the reasons behind the “anomalous” behavior of bootstrap current mentioned above. In that section, we also extend the alternative derivation of Helander et al. (2011) within the adjoint approach to the next order in collisionality. Although the obtained correction has no effect on the resulting Ware pinch (bootstrap) coefficient, it is useful for understanding a certain paradox contained in this solution. Collisionless asymptotic solutions are compared to numerical solutions for finite plasma collisionality in Section 2.3 where the cases with convergence of these solutions to the Shaing-Callen limit at low collisionality are demonstrated. In Section 3 we examine the effect of collisional boundary layers on the distribution function in the adjoint (Ware pinch) problem where they lead to the off-set of this function from the value of collisionless asymptotic of Helander et al. (2011) and respective off-set of the Ware-pinch coefficient. In particular, in Section 3.2 we introduce a simplified approach (propagator method) to describe this off-set in the leading order over collisionality with help of a set of Wiener-Hopf type integral equations. This set is infinite in case of irrational field lines, and becomes finite for closed field lines. In Section 3.3, with the help of this set, we derive the conditions on the equilibrium magnetic field required to avoid the leading order off-set. We obtain a simple expression for the distribution function off-set in case these conditions are weakly violated in Section 3.5 where we express these solutions via two discrete functions tabulated using the numerical solutions of two respective infinite integral equation sets resulting from the linearization of the original Wiener-Hopf type set. In Section 4 we examine the off-set of the distribution function and Ware pinch coefficient at irrational flux surfaces both, numerically, with help of NEO-2 code solutions at high order rational magnetic field lines approximating the irrational surface (Section 4.1), and using the analytical estimates of the asymptotic behavior of the off-set with decreasing plasma collisionality (Section 4.2). A related issue of bootstrap resonances is briefly discussed in Section 4.4. In Section 5 we study the effect of banana orbit precession (in particular, due to a finite radial electric field) on the off-set of Ware pinch coefficient and formulate there a simple bounce averaged approach for the account of this effect in computations of Ware pinch coefficient using the numerical solutions for the distribution function in the $1/\nu$ regime. A qualitative discussion of the off-set in the direct problem describing bootstrap coefficient is presented in Section 6 where we also examine the possibility of bootstrap effect at the magnetic axis. Finally, the results are summarized in Section 7 where some implications for reactor optimization are also discussed.
@@ 2.1 Adjoint mono-energetic problems on a closed field line
For the present analysis of bootstrap current convergence with plasma collisionality, a mono-energetic approach is sufficient where a Lorentz collision model and constant electrostatic potential within flux surfaces are assumed. The linear deviation of the distribution function $f$ from the local Maxwellian $f_{M}$ is presented as a superposition of thermodynamic forces $A_{k}$ as detailed in (Kernbichler et al., 2016),
@@
$H_{\varphi}^{\prime}=\frac{|\lambda|}{3B^{\varphi}}\left(3+\lambda^{2}\right)|\nabla r|k_{G}\rho_{L}$ (5)
will be integrated along the field line later to give bounce integrals $H_{j}$ in Eq. (26). Here, we use a field aligned coordinate system $(r,\vartheta_{0},\varphi)$ where $r$ is a flux surface label (effective radius) fixed by the condition $\langle|\nabla r|\rangle=1$, $\vartheta_{0}=\vartheta-\iota\varphi$ is a field line label, $\iota$ is the rotational transform and $\vartheta$ and $\varphi$ are poloidal and toroidal angle of periodic Boozer coordinates (Boozer, 1981; d’Haeseleer et al., 1991), respectively. Variables in the velocity space are parallel velocity sign $\sigma=v_{\parallel}/|v_{\parallel}|$ and two invariants of motion, $z=m_{\alpha}v^{2}/(2T_{\alpha})$ and $\eta=v_{\perp}^{2}/(v^{2}B)$ respectively being the normalized kinetic energy (playing a role of parameter) and perpendicular adiabatic invariant (magnetic moment). The other notation is pitch parameter $\lambda=v_{\parallel}/v=\sigma\sqrt{1-\eta B}$, the mean free path $l_{c}=v/(2\nu_{\perp})$ defined via deflection frequency $\nu_{\perp}$ (the same as $\nu$ in Ref. (Beidler et al., 2011)), radial guiding center velocity $v_{g}^{r}={\bf v}_{g}\cdot\nabla r$, Larmor radius $\rho_{L}=cm_{\alpha}v/(e_{\alpha}B)$ and the geodesic curvature given by
@@
with ${\bf h}={\bf B}/B$ and $\psi$ being the toroidal flux normalized by $2\pi$, and $B^{\varphi}$, $B_{\varphi}$ and $B_{\vartheta}$ being contra- and covariant components of the magnetic field in Boozer coordinates.
@@
For the present analysis, we solve the kinetic equation on the rational surface using a “representative” field line where necessary conditions valid for the irrational flux surface are fulfilled. The flux surface average (10) at the irrational surface corresponds then to the limit of the series of representative field lines closed after $N$ turns at respective rational surfaces, $r=r_{N}=M/N$, which converge to the irrational surface, $\lim\limits_{N\rightarrow\infty}\iota(r_{N})=\iota(r)$. The requested condition is Liouville’s theorem, which states that the integral over any closed surface of the normal component of the guiding center velocity multiplied by the phase space Jacobian is zero for fixed total energy and perpendicular adiabatic invariant used as phase space variables. For the magnetic surface with constant electrostatic potential, this means $\left\langle Bv_{g}^{r}/v_{\parallel}\right\rangle=0$, where the surface integration is performed over regions where $v_{\parallel}^{2}=v^{2}(1-\eta B)\geq 0$ keeping invariants $z$ and $\eta$ constant. The field line average form (10) of Liouville’s theorem results in
@@
for passing and trapped particles, respectively (see (4)). We will call these conditions “quasi-Liouville’s theorem”. In the devices with stellarator symmetry conditions (15) are satisfied for closed field lines passing through the magnetic field symmetry point $\varphi_{\text{sts}}$ what is obvious due to anti-symmetry of the geodesic curvature (and, respectively, of $v_{g}^{r}$ and $s_{(1)}$) with respect to this point, $s_{(1)}(2\varphi_{\text{sts}}-\varphi)=-s_{(1)}(\varphi)$. Restricting our analysis to such devices with a single global maximum per field period, the reference field line is then the one starting from the global maximum, which must be located at one of (at least two possible) symmetry points $\varphi_{0}\in\varphi_{\text{sts}}$.
@@ 2.2 Collisionless asymptotic solutions
where $\varphi_{\text{beg}}=\varphi_{j}^{-}(\eta)$ for trapped particles and $\varphi_{\text{beg}}=\varphi_{0}$ for passing, and we require that each of $g_{n}$ is continuous at the periodic boundary and at the turning points. Continuity at $\varphi=\varphi_{j}^{-}$ means that $\bar{g}_{n}$ is an even function of $\sigma$ in the trapped region, $\bar{g}_{n}=\bar{g}_{n}(\eta)$, while continuity at the periodic boundary in the passing region and at $\varphi=\varphi_{j}^{+}$ in the trapped region results in an equation for the integration constant $\bar{g}_{n-1}$ (solubility constraint for $g_{n}$). The leading order constraint for $g_{0}$ gives a bounce-averaged equation for $g_{-1}$
@@
where $\varphi_{\text{end}}=\varphi_{j}^{+}(\eta)$ for trapped particles and $\varphi_{\text{end}}=\varphi_{N}$ for passing, and the sum $\sum\limits_{\sigma=\pm 1}^{\text{trapped}}$ is taken for trapped particles only. Higher order constraints result in equations for integration constants $\bar{g}_{n}$,
@@
For the next order correction $g_{0}$, we notice that the first two terms in the r.h.s. of Eq. (19) are odd in $\sigma$. Therefore, in the trapped particle region, equation (23) for the integration constant (which can only be even there) is homogeneous, resulting in $\bar{g}_{0}=\text{const}$. Due to the continuity of the distribution function across the trapped-passing boundary, the integration constant in the passing region is the same. We can again absorb this constant into the equilibrium Maxwellian, as we have already done in the previous order when setting $g_{-1}=0$ in the passing region. Thus, substituting in (19) the derivative $\partial g_{-1}/\partial\eta$ via (27) and the source $s=s_{(1)}$ via (4) and (5) we get explicitly
@@
where we have exchanged the derivative over $\eta$ with integration in the first term using $H_{\varphi}^{\prime}(\varphi_{j}^{-}(\eta),\eta)=0$. It should be noted now that the integral over $\varphi^{\prime}$ is a discontinuous function of $\eta$ at the boundaries between classes $\eta=\eta_{c}$ where either $\varphi_{j}^{-}(\eta)$ or $\varphi_{j}^{+}(\eta)$ jumps. Respectively, the function $g_{0}$ contains a $\delta$-function $\delta(\eta-\eta_{c})$, which is required by particle conservation in the boundary layer. Namely, integrating Eq. (3) over $\eta$ across the boundary, we get
@@
$C_{\parallel}=C_{\parallel}(r)=\frac{1}{\rho_{L}}\sum_{\alpha^{\prime}}C_{\alpha^{\prime}}\rho_{L}^{\prime}=\frac{3c}{4B\rho_{L}}\frac{\partial p}{\partial r},\qquad p=\sum_{\alpha}p_{\alpha},$ (38)
$\rho_{L}^{\prime}$ denotes the Larmor radius of $\alpha^{\prime}$ species, and we used the condition $\bar{g}_{0}=0$ at the boundary $\eta=1/B$ when integrating by parts in the last expression (37). According to (35), the term $g_{0}-\bar{g}_{0}$ is a derivative which contributes only at the lower limit $\eta=0$ where using explicitly (5) one gets parallel equilibrium current density as
@@
which is a “true” rational surface condition (Solov’ev and Shafranov, 1970). Using the explicit form of $H^{\prime}_{\varphi}$ in (34) and taking the flux “surface average” (10) of $j_{\parallel}B$ thus eliminating the Pfirsch-Schlüter current which is driven solely by charge separation potential and, therefore, must satisfy $\langle j_{\text{PS}}B\rangle=0$, we finally obtain the bootstrap current density $j_{b}$ which scales with $B$ as
@@
The separate contribution of trapped particle region to $\lambda_{bB}$ is obtained by replacement of the lower integration limit in (37) from 0 to $\eta_{b}+o$,
@@
The factor (42) matches the result of Shaing and Callen, what is a natural consequence of using the same method. Namely, $\lambda_{bB}=(1-f_{c})\langle\tilde{G}_{b}\rangle/S$ where $S={\rm d}V/{\rm d}r$ is the flux surface area, fraction of circulating particles $f_{c}$ is given by Eq. (56) of Shaing and Callen (1983), and geometrical factor $\tilde{G}_{b}$ is given there by Eqs. (75b) and (60c). In case of negligible toroidal equilibrium current, $B_{\vartheta}=0$, the result of Boozer and Gardner (1990) is also approximately recovered, $\lambda_{bB}=\Delta_{0}B_{\varphi}{\rm d}r/{\rm d}\psi+O(\varepsilon_{\text{M}})$, where the quantity $\Delta_{0}$ is given by Eqs. (51), (53) and (54) of Boozer and Gardner (1990) ignoring small trapped particle contribution and other corrections linear in $\varepsilon_{M}$ (in case of a circular tokamak, $\lambda_{bB}\approx A\Delta_{0}$ where $A$ is the aspect ratio).
@@
Note that the trapped particle distribution function here is formally the same as the one given in the implicit form by Eq. (54) of Shaing and Callen (1983) omitting there the ripple plateau contribution absent in our case where we ignored banana precession (cross-field drift within flux surfaces) already in the starting equation (3). On the other hand, we obtained also the explicit expression for its odd part (35) which is demonstrative for the distribution of parallel equilibrium current in the velocity space. It can be seen that a significant part of this current (essentially of the Pfirsch-Schlüter current) flows in the boundary layers between trapped particle classes (but not in the main, trapped-passing boundary layer). The origin of these localized currents is the same as the origin of Pfirsch-Schlüter current in all devices, i.e. they remove charge separation introduced by the radial particle drift. However, in contrast to tokamaks (and tokamak-like part of the current in stellarators) where compensation currents are produced in a long mean free path regime within a single turn of particle bounce motion due to the finite radial displacement during this time, localized currents serve to compensate charge separation by finite bounce-averaged drift, which accumulates on much longer collisional detrapping time. Interpreting the source term $s_{(1)}$, Eq. (4), in the regions where it is positive as a source of “particles” and where it is negative as a source of “anti-particles” whose total amount generated at the flux surface is the same as the amount of “particles” being a consequence of quasi-Liouville’s theorem (15), we see that particles (or anti-particles) accumulated in the local ripple wells due to the finite bounce averaged drift, $H_{j}\neq 0$, can only leave the wells due to collisional scattering flux through the class boundaries (see Fig. 1). Since local collisional flux density generally needs not to be continuous at this boundary, $\partial g_{-1}(\eta-o)/\partial\eta\neq\partial g_{-1}(\eta+o)/\partial\eta$, a significant amount of particles (anti-particles) is re-distributed through the boundary layers, what is manifested by the $\delta$-like behavior of $g_{0}$, Eq. (35), being, up to a factor, a parallel flow density in velocity space.
@@
The asymptotic series expansion (17) can be continued to the next order, leading to the correction $g_{1}$. This correction, however, has a non-integrable singularity at class transition boundaries, $g_{1}\propto(\eta-\eta_{c})^{-2}$ (in contrast to integrable singularity of $g_{0}\propto\log|\eta-\eta_{c}|$), which is a consequence of finite geodesic curvature at respective local field maxima. Therefore, contributions of different order corrections to the collisional flux density in the matching conditions (31) become comparable at finite collisionality for $|\eta-\eta_{c}|=\delta\eta\propto l_{c}^{-1/2}\propto\nu^{1/2}$ where $\partial g_{0}(\eta)/\partial\eta\sim\delta\eta^{-1}$ and $\partial g_{1}(\eta)/\partial\eta\sim l_{c}^{-1}\delta\eta^{-3}$. In other words, the expansion (17) breaks down at the edge of the boundary layer of width $\delta\eta$. Respectively, matching interval $o$ in Eq. (28), which is an infinitesimal for vanishing collisionality, should satisfy $\delta\eta\ll o\ll\eta$ for finite collisionality, which means that the error of the asymptotic solution scales with $\sqrt{\nu}$ rather than with $\nu$. Since particle – anti-particle re-distribution flux manifested by $\delta$-functions in asymptotic expression (35) for $g_{0}$ is independent of collisionality, we can estimate the odd part of the distribution function which carries this flux in the class-transition boundary layers as $g^{\text{odd}}\propto\delta\eta^{-1}\propto l_{c}^{1/2}\propto\nu^{-1/2}$.
As shown below, the presence of these strong parallel flows in class transition boundary layers is actually responsible for the “anomalous” behavior of bootstrap current at finite collisionality.